# Federated Anonymous Blocklisting across Service Providers and its Application to Group Messaging

### v1

@@ Abstract
Instant messaging has become one of the most used methods of communication online, which has attracted significant attention to its underlying cryptographic protocols and security guarantees. Techniques to increase privacy such as End-to-End Encryption and pseudonyms have been introduced. However, online spaces such as messaging groups still require moderation to prevent misbehaving users from participating in them, particularly in anonymous contexts.. In Anonymous Blocklisting (AB) schemes, users must prove during authentication that none of their previous pseudonyms has been blocked, preventing misbehaving users from creating new pseudonyms. In this work we propose an alternative Federated Anonymous Blocklisting (FAB) in which the centralised Service Provider is replaced by small distributed Realms, each with its own blocklist. Realms can establish trust relationships between each other, such that when users authenticate to a realm, they must prove that they are not banned in any of its trusted realms. We provide an implementation of our proposed scheme; unlike existing AB constructions, the performance of ours does not depend on the current size of the blocklist nor requires processing new additions to the blocklist. We also demonstrate its applicability to real-world messaging groups by integrating our FAB scheme into the Messaging Layer Security protocol.
@@ 1 Introduction
In recent years, instant messaging has become the main method of communication over the internet [messaging-stats], which has sparked interest from the scientific community to analyse and improve their underlying security protocols. End-to-End Encryption (E2EE) [e2ee] ensures that messages can only be decrypted by their intended recipients and not by any intermediate servers. Tools for anonymity have also been expanded: Signal’s sealed sender [signal-sender] allows users to send messages such that the routing servers do not learn the identity of the sender. These efforts to expand the security guarantees of instant messaging have been incorporated into widely used applications, such as WhatsApp due to its adoption of the Signal protocol [signal]. The Messaging Layer Security (MLS) protocol [mls], recently standardised as RFC 9420, efficiently expands E2EE to the context of group messaging.
@@
At the same time, the service providers that own said messaging applications may desire to moderate the content shared through them, whether due to legal obligations or to prevent harassment or other forms of misconduct. Moderation becomes more difficult as the users’ security guarantees increase: indeed, it is harder for service sroviders to moderate spaces in which users are anonymous or messages are deniable —that is, it is impossible to prove that an specific user sent a given message. One solution is message franking[franking], which allows participants to prove to the service provider that some other user has acted maliciously, selectively breaking E2EE —and for Sealed Sender, also anonymity [franking-sealed]— to introduce moderation.
@@
On the other hand, Anonymous Blocklisting (AB) schemes propose an alternative moderation tool that does not compromise the anonymity of users. In these schemes users interact with service providers anonymously through the use of pseudonyms derived from their identity. In order to authenticate, users provide a (zero-knowledge) proof that none of their previous pseudonyms has been blocked. While undoubtedly useful, the performance of AB schemes often scales poorly with the number of blocked users [blac, snarkblock]. Furthermore, these schemes require users to process every new insertion to the blocklist —even if they are not affected [alpaca]. As target service providers are estimated to block thousands of users per day, the efficiency of current AB schemes quickly degrades.
@@
Both AB and message franking share in common that in order to introduce moderation, they require a powerful service provider that is able to arbitrarily enforce blocks across its user base and potentially access the contents of conversations between users. This assumption does not hold for decentralised environments in which each association of users (such as instant messaging groups) may decide their own policies for blocking users. For example, the so-called fediverse allows every instance to define their policies, which are internally enforced by their administrators [fediverse].
@@
In this work, we adapt the Anonymous Blocklisting framework to this environment: instead of assuming a centralised service provider with a single blocklist, our model considers a large number of small Realms, independent logical domains, each maintaining its own blocklist. Users employ different pseudonyms in each Realm, such that their activity across Realms cannot be linked. However, Realms can decide to enforce the blocklist of a different Realm, such that users attempting to authenticate to a Realm must also prove that they are not blocked in any of its trusted Realms. The resulting scheme is called Federated Anonymous Blocklisting (FAB).
@@
We provide a construction that fulfils the requirements of a FAB scheme and develop a formal security analysis to prove the security properties of Unlinkability, Blocklistability and Unframeability defined for AB schemes [ab_formal]. Our construction introduces significant improvements in efficiency over existing AB schemes. Crucially, its performance scales logarithmically with the maximum size of the blocklist. While in other AB schemes users need to process every new insertion to the blocklist, even if they are not affected, our construction completely avoids that cost. However, users need to generate a proof for every realm in which the target realm trusts, introducing a different linear scaling to the performance.
@@ 2 Related Work
Anonymous Blocklisting schemes allows users to authenticate to Service Providers using unlinkable pseudonyms. Service Providers also hold a blocklist of pseudonyms, such that blocked users are unable to successfully authenticate again [ab_formal]. The first AB scheme was presented in [blac]; in this work, authentication costs scaled linearly with the blocklist size. Similarly, the authors of [perea] present a similar scheme that limits the analysis to the most recent authentication attempts by the user.
To improve efficiency, other AB works introduce techniques that allows users to reuse proofs. The SNARKBlock scheme [snarkblock] uses aggregable zk-SNARKs to divide the blocklist into chunks such that chunk proofs could be reused. While a significant improvement, the cost of aggregating the chunk proofs still increases linearly with block size. The recent ALPACA [alpaca] instead employs Incremental Verifiable Computation [ivc] to achieve performance costs independent of blocklist size. Both of these works require users to perform an expensive offline synchronisation whenever new elements are inserted into the blocklist.
Federated identity systems are composed of multiple Service Providers and Identity Providers with which users interact in order to authenticate. To provide unlinkability across various Service Providers, users employ different pseudonyms in each of them generated from the user’s identity and a value provided by the Service Provider, called scoped Pseudonyms. In [attr_pseudo], the pseudonyms are generated from attribute credentials. The scheme defined in [bison] also prevents Identity Providers from linking the user’s activity by employing blind evaluation. These works allow users to be traced in a specific scope but unlinkable between scopes. However, there is no mechanism for preventing users blocked in one scope to authenticate in other scopes.
There have been recent efforts to increase anonymity and privacy in messaging groups while maintaining access control. Usually, open standards for end-to-end encryption are employed for these schemes, such as Signal [signal] and Messaging Layer Security [mls]. The IETF drafts [mimi, mimimi] introduce metadata minimalisation groups in which the member’s credentials are encrypted and are only accessible to other group members. In [aa_cgka], a scheme for anonymously authenticating in messaging groups using Attribute-Based Credentials and fresh key pairs is defined. The authors of [orca] apply Anonymous Blocklisting to protect recipients of sender-anonymous messages. Message franking [franking, franking-sealed] is also employed for moderation in end-to-end encrypted messaging. However, message franking breaks the confidentiality of end-to-end encryption by allowing moderators to access the contents of reported messages.
@@ 3 Background
Completeness. The probability
$\operatorname{Pr}\left[\begin{array}[]{c}((pk,vk),td)\leftarrow\mathsf{Gen}(R)\\ \pi\leftarrow\mathsf{Prove}(pk,x,w)\\ (x,w)\in R\\ \hline\cr\mathsf{Verify}(vk,x,\pi)=Reject\end{array}\right]$
is negligible. Intuitively, this means that an honest prover is able to convince a verifier that $(x,w)\in R$.
@@ 3.2 Negative Accumulators
A Cryptographic Accumulator is a data structure that concisely represents a set of values [acc]. Alongside its benefits in space, accumulators also provide computationally efficient methods for creating and verifying (non-)membership proofs that ensure that a specific element is (or is not) in the set of values. While multiple families of accumulators exist [acc_survey, acc_survey_2, acc_survey_3], we are only interested in negative accumulators [acc_universal]. These accumulators allow for the creation on non-membership proofs, which can be used to prove that a certain element has not been inserted into it. We also require a dynamic accumulator that allows insertions and deletions to its set of values. Formally, a negative accumulator is defined by the following operations [oblivious_acc]:
@@
Negative accumulators possess the following properties:
Correctness. An honestly-generated non-membership proof for an element not present in the accumulator will be accepted by an honest verifier with overwhelming probability.
Soundness. For every efficient adversary $A$, the probability
@@ 4.1 System model
The system model of the FAB scheme is represented in Figure 1. The following entities participate in a FAB scheme:
  • 1.
    Users ($U$): possess a private identity $x$. Their objective is to successfully authenticate to realms using a pseudonym.
  • 2.
    Identity Provider ($IP$): Creates credentials for users in which they sign their identity. They are trusted by the other entities in the scheme.
  • 3.
    Realms ($R$): abstract entities that verify the users’ identity. Each realm possesses a blocklist to which user pseudonyms can be inserted into at the realm’s discretion.
We assume the Identity Provider is able to ensure Sybil resistance, that is, users cannot possess multiple credentials referencing different identities. Furthermore, whereas we assume that every entity trusts on $IP$’s signature, we require that it never learns the identity of users.
@@
For practical purposes we define an auxiliary entity called Realm Directory, which users can query to obtain public information about any realm. In practice, this Directory could simply forward the request to a maintainer of the realm, or consist on a server that periodically receives updates from realms.
@@ 4.2 Comparison with Anonymous Blocklisting schemes
In order to contextualise our work we highlight its similarities and differences with other Anonymous Blocklisting schemes [ab_formal]. Both schemes operate in the same context: Users employ pseudonyms to interact with Service Providers, which can discretionarily block pseudonyms. AB’s Service Providers are renamed as Realms in this work to emphasise the coexistence of an indefinite amount of them with their own context and members. Blocks in AB schemes are meant to prohibit the user from authenticating again in the same Service Provider; this also applies to our FAB scheme, but our main contribution is that it also prevents the user from authenticating in other realms.
@@ A FAB scheme consists of the following algorithms:
    $st=(s,acc)\leftarrow\mathsf{Create}(1^{\lambda})$: Creates a realm with state $st$, which is composed of a seed $s$ and an accumulator accacc.
@@
Figure 2 shows an example execution flow of a FAB scheme. First, an user $U$ registers their identity to the Identity Provider and receives a credential signed by it. In order to authenticate to target realm $R$, $U$ must first obtain $R$’s state $st_{R}$ and the state of all of its trusted realms $\{st_{i}\}_{i\in I}$. Once $U$ receives this information from the Realm Directory, the $\mathsf{Auth}$ algorithm is executed to create a pseudonym $ps^{R}_{U}$, a proof of non-membership to $R$’s blocklist $\overline{w^{R}_{U}}$ and a proof $\pi$. When $R$ receives the authentication attempt, it first retrieves the state of its trusted realms in order to obtain the latest updates. Then, it executes the $\mathsf{Verify}$ algorithm to check the validity of the proof.
Figure 2: Authentication flow of the FAB scheme. The User $U$ first registers its identity with the Identity Provider and then obtains the list of realms trusted by $R$. $U$ then executes $\mathsf{Auth}$ to prove they are not banned, which is later verified by $R$.
@@ 4.4 Security
Our security analysis follows the standard Anonymous Blocklisting framework of [blac, snarkblock]: in particular, we also employ the security notions of Blocklistability, Unlinkability and Non-Frameability. However, we provide custom definitions of said properties to account for the differences in our system model: while the referenced works define a more general version of Anonymous Blocklisting, our scenario involves multiple realms each with its own blocklist and trust relations between each other.
We follow the query-based approach of [alpaca]. The security properties are defined as a game between a Challenger and an Adversary. Each security game has its own winning condition and may incorporate checks to prevent trivial wins by the adversary. Algorithm 1 shows the shared queries between all security games —with a slight variation in the $Q_{Block}$ query for the Non-Frameability and Unlinkability games. They are mostly related to administrative tasks such as creating realms, registering users and blocking and unblocking pseudonyms.
Algorithm 1 Shared queries and auxiliary functions through all security games. Blue-highlighted text in the Block query only applies to the Non-Frameability game. Red-highlighted text applies only to the Unlinkability game.
@@
Algorithm 2 Blocklistability Security game.
@@
  $C,R\leftarrow\varnothing$
  $creds,realms\leftarrow 0$
@@
Algorithm 3 Unlinkability Security game.
@@
  $A,R,B\leftarrow\varnothing$
  $creds,realms\leftarrow 0$
@@
Algorithm 4 Non-Frameability Security game.
@@
  $C,R,B\leftarrow\varnothing$
  $creds,realms\leftarrow 0$
@@ 4.4.1 Blocklistability
The Blocklistability security property states that a successful authentication can only be generated by a registered user that is not blocklisted neither in the target realm nor in any of its trusted realms. Thus, this property also considers authentication attempts by unregistered users, which is sometimes referred as mis-authentication resistance [blac].
The security game is shown in Algorithm 2, and it unfolds as follows: the Challenger initialises the environment and allows the adversary to execute arbitrary queries. Eventually, the adversary outputs a target realm, a set of trusted realms, and an authentication attempt. The adversary wins if said authentication attempt is valid but all created users are banned on either the target realm or on at least one of the trusted realms.
@@ 4.4.2 Unlinkability
The Unlinkability property ensures that it is impossible to identify if two pseudonyms in different realms correspond to the same user without knowing the original credential.
The security game is shown in Algorithm3, and it unfolds as follows: the Challenger creates two users with credentials $cred_{0},cred_{1}$, and the adversary is allowed to execute arbitrary queries. Eventually the adversary queries $Q_{Chal}$ with references to a realm and a set of trusted realms. The Challenger then authenticates $cred_{b}$ in said realm. The adversary wins if they successfully guess $b$. The $Q_{Auth}$ query allows the adversary to obtain honestly-generated authentication attempts from the challenge credentials, as the adversary cannot generate them by themselves. Trivial attacks are prevented by the following checks: (1) querying $Q_{Chal}$ for a realm in which either user would not be able to generate a valid proof or (2) the adversary already knows the pseudonym of a user, (3) querying $Q_{Auth}$ for the challenge realm or (4) blocking the challenge pseudonym, which would provoke future $Q_{Auth}$ queries to fail only for the challenge user. The latter check is highlighted in green in Algorithm 1.

### v2

@@ Abstract
Instant messaging has become one of the most used methods of communication online, which has attracted significant attention to its underlying cryptographic protocols and security guarantees. Techniques to increase privacy such as End-to-End Encryption and pseudonyms have been introduced. However, online spaces such as messaging groups still require moderation to prevent misbehaving users from participating in them, particularly in anonymous contexts.. In Anonymous Blocklisting (AB) schemes, users must prove during authentication that none of their previous pseudonyms has been blocked, preventing misbehaving users from creating new pseudonyms. In this work we propose an alternative Federated Anonymous Blocklisting (FAB) in which the centralised Service Provider is replaced by small distributed Realms, each with its own blocklist. Realms can establish trust relationships between each other, such that when users authenticate to a realm, they must prove that they are not blocked in any of its trusted realms. We provide an implementation of our proposed scheme; unlike existing AB constructions, the performance of ours does not depend on the current size of the blocklist nor requires processing new additions to the blocklist. We also demonstrate its applicability to real-world messaging groups by integrating our FAB scheme into the Messaging Layer Security protocol.
keywords:
Anonymous Blocklisting, Messaging groups, zk-SNARKs, MLS
@@ 1 Introduction
In recent years, instant messaging has become the main method of communication over the internet [33], which has sparked interest from the scientific community to analyse and improve their underlying security protocols. End-to-End Encryption (E2EE) [13] ensures that messages can only be decrypted by their intended recipients and not by any intermediate servers. Tools for anonymity have also been expanded: Signal’s sealed sender [31] allows users to send messages such that the routing servers do not learn the identity of the sender. These efforts to expand the security guarantees of instant messaging have been incorporated into widely used applications, such as WhatsApp due to its adoption of the Signal protocol [1]. The Messaging Layer Security (MLS) protocol [6], recently standardised as RFC 9420, efficiently expands E2EE to group messaging.
@@
At the same time, the service providers that own said messaging applications may desire to moderate the content shared through them, whether due to legal obligations or to prevent harassment or other forms of misconduct. Moderation becomes more difficult as the users’ security guarantees increase: indeed, it is harder for service sroviders to moderate spaces in which users are anonymous or messages are deniable —that is, it is impossible to prove that an specific user sent a given message. One solution is message franking[12], which allows participants to prove to the service provider that some other user has acted maliciously, selectively breaking E2EE —and for Sealed Sender, also anonymity [17]— to introduce moderation.
@@
On the other hand, Anonymous Blocklisting (AB) schemes propose an alternative moderation tool that does not compromise the anonymity of users. In these schemes users interact with service providers anonymously through the use of pseudonyms derived from their identity. In order to authenticate, users provide a (zero-knowledge) proof that none of their previous pseudonyms has been blocked. While undoubtedly useful, the performance of AB schemes often scales poorly with the number of blocked users [35, 28]. Furthermore, these schemes require users to process every new insertion to the blocklist —even if they are not affected [20]. As target service providers are estimated to block thousands of users per day, the efficiency of current AB schemes quickly degrades.
@@
Both AB and message franking share in common that in order to introduce moderation, they require a powerful service provider that is able to arbitrarily enforce blocks across its user base and potentially access the contents of conversations between users. This assumption does not hold for decentralised environments in which each association of users (such as instant messaging groups) may decide their own policies for blocking users. For example, the so-called fediverse allows every instance to define their policies, which are internally enforced by their administrators [2].
@@
In this work, we adapt the Anonymous Blocklisting framework to this environment: instead of assuming a centralised service provider with a single blocklist, our model considers a large number of Realms, independent logical domains, each maintaining its own blocklist. Users employ different pseudonyms in each Realm, such that their activity across Realms cannot be linked. Realms can decide to enforce the blocklist of a different Realm, such that users attempting to authenticate to a Realm must also prove that they are not blocked in any of its trusted Realms. The resulting scheme is called Federated Anonymous Blocklisting (FAB).
@@
We provide a construction that fulfils the requirements of a FAB scheme and develop a formal security analysis to prove the security properties of Unlinkability, Blocklistability and Unframeability defined for AB schemes [15]. Our construction introduces significant improvements in efficiency over existing AB schemes. Crucially, its performance scales logarithmically with the maximum size of the blocklist. While in other AB schemes users need to process every new insertion to the blocklist, even if they are not affected, our construction completely avoids that cost. However, users need to generate a proof for every realm in which the target realm trusts, introducing a different linear scaling to the performance.
@@ 2 Related Work
Anonymous Blocklisting schemes allows users to authenticate to Service Providers using unlinkable pseudonyms. Service Providers also hold a blocklist of pseudonyms, such that blocked users are unable to successfully authenticate again [15]. The first AB scheme was presented in [35]; in this work, authentication costs scaled linearly with the blocklist size. Similarly, the authors of [34] present a similar scheme that limits the analysis to the most recent authentication attempts by the user.
To improve efficiency, other AB works introduce techniques that allows users to reuse proofs. The SNARKBlock scheme [28] uses aggregable zk-SNARKs to divide the blocklist into chunks such that chunk proofs could be reused. While a significant improvement, the cost of aggregating the chunk proofs still increases linearly with block size. The recent ALPACA [20] instead employs Incremental Verifiable Computation [22] to achieve performance costs independent of blocklist size. Both of these works require users to perform an expensive offline synchronisation whenever new elements are inserted into the blocklist.
Federated identity systems are composed of multiple Service Providers and Identity Providers with which users interact in order to authenticate. To provide unlinkability across various Service Providers, users employ different pseudonyms in each of them generated from the user’s identity and a value provided by the Service Provider, called scoped Pseudonyms. In [9], the pseudonyms are generated from attribute credentials. The scheme defined in [14] also prevents Identity Providers from linking the user’s activity by employing blind evaluation. These works allow users to be traced in a specific scope but unlinkable between scopes. However, there is no mechanism for preventing users blocked in one scope to authenticate in other scopes.
There have been recent efforts to increase anonymity and privacy in messaging groups while maintaining access control. Usually, open standards for E2EE are employed for these schemes, such as Signal [1] and Messaging Layer Security [6]. The IETF drafts [7, 21] introduce metadata minimalisation groups in which the member’s credentials are encrypted and are only accessible to other group members. In [32], a scheme for anonymously authenticating in messaging groups using Attribute-Based Credentials and fresh key pairs is defined. The authors of [36] apply Anonymous Blocklisting to protect recipients of sender-anonymous messages. Message franking [12, 17] is also employed for moderation in E2EE messaging to allow moderators to identify the original source of a message.
@@ 3.2 Negative Accumulators
A Cryptographic Accumulator is a data structure that concisely represents a set of values [8]. Alongside its benefits in space, accumulators also provide computationally efficient methods for creating and verifying (non-)membership proofs that ensure that a specific element is (or is not) in the set of values. While multiple families of accumulators exist [25, 24, 27], we are only interested in negative accumulators [23]. These accumulators allow for the creation on non-membership proofs, which can be used to prove that a certain element has not been inserted into it. We also require a dynamic accumulator that allows insertions and deletions to its set of values. Formally, a negative accumulator is defined by the following operations [5]:
@@
Negative accumulators possess the Soundness property, which states that for every efficient adversary $A$, the probability
@@ 4.1 System model
The system model of the FAB scheme is represented in Figure 1. The following entities participate in a FAB scheme: Users ($U$) possess a private identity $x$, and their objective is to successfully authenticate to realms using a pseudonym. Realms ($R$) are abstract entities that verify the users’ identity, and each realm possesses a blocklist to which user pseudonyms can be inserted into at the realm’s discretion.
Additionally, the system requires an Identity Provider ($IP$) to create credentials for users in which they sign their identity. $IP$’s signature is trusted by the other entities in the scheme, but it is not trusted to learn the users’ identity. We assume the $IP$ is able to ensure Sybil resistance, that is, users cannot possess multiple credentials referencing different identities.
@@
For practical purposes we define an auxiliary entity called Realm Directory, which users can query to obtain public information about any realm. In practice, this Directory could forward the request to a maintainer of the realm, or consist on a server that periodically receives updates from realms.
@@ 4.2 Comparison with Anonymous Blocklisting schemes
In order to contextualise our work we highlight its similarities and differences with other Anonymous Blocklisting schemes [15]. Both schemes operate in the same context: Users employ pseudonyms to interact with Service Providers, which can discretionarily block pseudonyms. AB’s Service Providers are renamed as Realms in this work to emphasise the coexistence of an indefinite amount of them with their own context and members. Blocks in AB schemes are meant to prohibit the user from authenticating again in the same Service Provider; this also applies to our FAB scheme, but our main contribution is that it also prevents the user from authenticating in other realms.
@@ A FAB scheme consists of the following algorithms:
    $st\leftarrow\mathsf{Create}(1^{\lambda})$: Creates a realm with state $st$.
@@
Figure 2 shows an example execution flow of a FAB scheme. First, an user $U$ registers their identity to $IP$ and receives a signed credential. In order to authenticate to target realm $R$, $U$ must first obtain from the Realm Directory $R$’s state $st_{R}$ and the state of all of its trusted realms $\{st_{i}\}_{i\in I}$. Then, $U$ executes the $\mathsf{Auth}$ algorithm to create a pseudonym $ps^{R}_{U}$, a proof of non-membership to $R$’s blocklist $\overline{w^{R}_{U}}$ and a proof $\pi$. When $R$ receives the authentication attempt, it first retrieves the state of its trusted realms to obtain the latest updates. Finally, it executes the $\mathsf{Verify}$ algorithm validate $\pi$.
Figure 2: Authentication flow of the FAB scheme. The User $U$ first registers its identity with the Identity Provider and then obtains the list of realms trusted by $R$. $U$ then executes $\mathsf{Auth}$ to prove they are not blocked, which is later verified by $R$.
@@ 4.4 Security
Our security analysis follows the standard Anonymous Blocklisting framework of [35, 28]: we also employ the security notions of Blocklistability, Unlinkability and Non-Frameability. However, we provide custom definitions of said properties to account for the differences in our system model.
We follow the query-based approach of [20]. The security properties are defined as a game between a Challenger and an Adversary. Each security game has its own winning condition and may incorporate checks to prevent trivial wins by the adversary. Figure 3 shows the shared queries between all security games —with a slight variation in the $Q_{Block}$ query for the Non-Frameability and Unlinkability games. They are mostly related to administrative tasks such as creating realms, registering users and blocking and unblocking pseudonyms.
@@
Figure 3: Shared queries and auxiliary functions through all security games. Blue-highlighted text in the Block query only applies to the Non-Frameability game. Red-highlighted text applies only to the Unlinkability game.
@@
  $C,R\leftarrow\varnothing;creds,realms\leftarrow 0$
@@
Figure 4: Blocklistability Security game.
@@
  $A,R,B\leftarrow\varnothing;creds,realms\leftarrow 0$
@@
Figure 5: Unlinkability Security game.
@@
  $C,R,B\leftarrow\varnothing;creds,realms\leftarrow 0$
@@ 4.4.1 Blocklistability
The Blocklistability property states that a successful authentication can only be generated by a registered user that is not blocklisted neither in the target realm nor in any of its trusted realms. Thus, this property also considers authentication attempts by unregistered users.
The security game is shown in Figure 4, and it unfolds as follows: the Challenger initialises the environment and allows the adversary to execute arbitrary queries. Eventually, the adversary outputs a target realm, a set of trusted realms, and an authentication attempt. The adversary wins if said authentication attempt is valid but all created users are blocked on either the target realm or on at least one of the trusted realms.
@@ 4.4.2 Unlinkability
The Unlinkability property ensures that it is impossible to identify if two pseudonyms in different realms correspond to the same user without knowing the user’s original credential.
The security game is shown in Figure5, and it unfolds as follows: the Challenger creates two users with credentials $cred_{0},cred_{1}$, and the adversary is allowed to execute arbitrary queries. Eventually the adversary queries $Q_{Chal}$ with references to a realm and a set of trusted realms. The Challenger then authenticates $cred_{b}$ in said realm. The adversary wins if they successfully guess $b$. The $Q_{Auth}$ query allows the adversary to obtain honestly-generated authentication attempts from the challenge credentials, as the adversary cannot generate them by themselves. Trivial attacks are prevented by the following checks: (1) querying $Q_{Chal}$ for a realm in which either user would not be able to generate a valid proof or (2) the adversary already knows the pseudonym of a user, (3) querying $Q_{Auth}$ for the challenge realm or (4) blocking the challenge pseudonym, which would provoke future $Q_{Auth}$ queries to fail only for the challenge user. The latter check is highlighted in green in Figure 3.